# On the modeling of irreversibility by relaxator Liouville dynamics

### v1

@@ IV Kinetic equations from relaxator Liouville dynamics
An alternative way to the BBGKY route to arrive at kinetic equations uses so called nonequilibrium Green functions (NEGF) (invented by Schwinger, Feynman and Vernon, Kadanoff and Baym, Keldysh Neqinventors and others in the 1960’s). These are nonequilibrium expectation values for one quantum object (particle) to propagate from one event $1$ to another event $2$. In the NEGF approach, the idea of a self-energy $\Sigma$ is exploited. It takes the many particle influences on the one-particle propagator into account on its way from $1$ to $2$. For an introduction to NEGF see VLeuwen2006 , Chap. 13 in Bonitz2016 and Datta . Usually, in the NEGF approach, the self-energy is constructed by approximate methods to incorporate as much of the interaction between many particles as possible or needed. The self-energy can then show finite broadened levels shifted from stationary one-particle levels.
The route to kinetic equations we take with relaxator Liouville dynamics is close in spirit to the NEGF method, but the construction of the self-energy is formally exact owing to a projector method. Our procedure has the advantage to immediately leading us to a closed equation for the one-particle density operator via a one-particle propagator. It makes the origin of irreversible behavior apparent by the relaxator appearing in this equation. The corresponding spectral shift and relaxator in our approach correspond to the self-energy in the NEGF approach.
@@ V Linear response for relaxator Liouville dynamics
Now, we expand this equation in powers of $\delta L{F}_{\tilde{\omega}}$ and estimate the regime of reliable linear response by requiring that the linear contribution to $\rho(z)$ is smaller than the contribution without the field. Since $(z-L(z))^{-1}$ scales with the system’s spectral exploration time $t_{s}$ and the coupling Liouville scales with a characteristic rate as the inverse of a field induced transition time $t_{\rm\delta F_{\omega}}$, the condition for the linear response theory to apply reads
$t_{s}\ll t_{\rm\delta F_{\omega}}\,.$ (97)
@@ VI Weak coupling approximation
Here $N(\Omega)$ is the average equilibrium number of modes. However, there is a problematic property of the linear coupling to independent bath modes: the correlation function $\gamma(\Omega)$ has no characteristic time scale involved besides the time scale inherent in $\rho_{\rm env}$ (e.g. the temperature time scale $T^{-1}$ for environmental equilibrium) and the time scale inherent in the density of levels $\nu(|\Omega|)$ (which is the environmental inverse level spacing $t_{e}(|\Omega|)$). The characteristic time scale of an environmental correlation time $\tau_{e}$ is missing as an ingredient in a bath of independent modes and can only be introduced phenomenologically. Typically, for independent modes in the limit of large volume the density of states becomes a power low $\sim\Omega^{p}$ with $p\geq 1$ such that $\gamma(0)\to 0$ (due to $\lim_{\omega\to 0}t_{e}(\omega)=\infty$) in the limit of continuous environmental spectrum. However, for a realistic long time correlation function with decay on a scale $\tau_{e}$ one should find $\gamma(0)\propto\tau_{e}\not=0$. Therefore, one shouldn’t take the correlation function of (144) literally, but as a means of parametrization to be adjusted phenomenologically or to be improved by couplings which allow to introduce characteristic correlation time scales of the environment. For a deeper discussion of limitations of the linear coupling to independent bath modes we refer to AlickiQDS and Horn2009 . In the following we will not rely on the linear coupling expression (144) but stick to the general properties of $\gamma(\Omega)$ for a bath in canonical equilibrium.
@@ VII Markov approximation: semi-group dynamics
as a semi-group-generator, because we want to be sure that its stationary limit is as close as possible with that of the original relaxator Liouville. The condition for the absence of memory (157) is quite restrictive and often not fulfilled in a realistic system when times between $\delta t_{s}$ and $\tau_{r}$ are addressed. One may ask why the semi-group approach that assumes the absence of memory (and absence of initial correlations) is still quite successful in describing several features of open systems. We think the answer can be given by looking at the solution (34) of the relaxator Liouville dynamics in its spectral resolution. The solution (38) shows a superposition of damped oscillations which could as well be described as the solution of a frequency independent Liouville $L$ with its eigenfrequencies adopted to the effective eigenfrequencies $z_{k}$ (37) of the full $L(z)$. In particular, when fast oscillatory behavior is averaged out, the Markov approximation, still capturing relaxation, becomes better.
@@
VIII Summary
We developed the relaxator Liouville dynamics (Eqs. (4),(16) - (19)) for relevant degrees of freedom in a macroscopic system from a microscopic von Neumann equation for an isolated total system of all degrees of freedom. The irrelevant degrees of freedom act as an environment on the relevant degrees of freedom. The relaxator $\Gamma(\omega)$ indicates the onset of irreversibility. It can be expressed by a Hamiltonian modeling of system environment coupling (Eq. (70). For a weak coupling the relaxator Liouville is determined by environmental correlation functions and hopping operators in the system (Eq. (122)). The relaxator Liouville dynamics has generically a unique stationary state (Eq. (48)) which is reached after a characteristic relaxation time that can be calculated from the relaxator Liouville’s effective eigenvalues (Eq. (39)). For equilibrium conditions the system reaches canonical thermal equilibrium with a temperature adopted to the environment.
@@ VIII Summary
For situations of current carrying stationary states in the presence of perturbing external fields we derived linear response expressions for dynamic susceptibilities (Eq. (102)). The derived susceptibilities coincide with Kubo’s susceptibilities in a regime where the energy eigenstates of the unperturbed system have undetectable large lifetimes and the probing frequency is much larger than the typical level spacing. Despite the practically infinite lifetimes of energy eigenstates, the response can show an collective irreversible character as the spectrum of the system becomes dense relative to the frequency. The corresponding time scale of irreversibility, a decay time of correlations within the system, can be extracted from the susceptibility (Eq. (117)).
@@
As the source of irreversibility we have identified time scale separations. For a system in contact with a surrounding bath the time to explore the spectral content of the bath is much larger than the time to explore the spectral content of the system. For a macroscopic system in general the dynamics of the so-called relevant variables becomes irreversible when the time to explore its spectral content is much smaller than the time to explore the spectral content of the complementary dynamics. In transport situations with driving fields, irreversibility and dissipation of energy emerges when they are macroscopic or subject to macroscopic contacts such that the level spacing of the system cannot be resolved with the frequency of the applied field. In other words: the time to explore the system’s spectral content is much larger as any time we can afford to measure the response of the system to applied fields.
@@
Acknowledgements The authors acknowledge valuable comments during the process of writing from Victor Albert (2018), Heinz-Peter Breuer (2018) , John Chalker (2024), Rochus Klesse (2017), Dimitry Polyakov (2025), János Polonyi (2017, 2025) and Peter Wölfle (2025).
@@
Keywords irreversibility, non-equilibrium, relaxation, quantum Master, non-Markovian, kinetic equations, linear response

### v2

@@ IV Kinetic equations from relaxator Liouville dynamics
An alternative way to the BBGKY route to arrive at kinetic equations uses so called nonequilibrium Green functions (NEGF) (invented by Schwinger, Feynman and Vernon, Kadanoff and Baym, Keldysh Neqinventors and others in the 1960’s). These are nonequilibrium expectation values for one quantum object (particle) to propagate from one event $1$ to another event $2$. In the NEGF approach, the idea of a self-energy $\Sigma$ is exploited. It takes the many particle influences on the one-particle propagator into account on its way from $1$ to $2$. For an introduction to NEGF see VLeuwen2006 , Chap. 13 in Bonitz2016 and Datta . Usually, in the NEGF approach, the self-energy is constructed by approximate methods to incorporate as much of the interaction between many particles as possible or needed. The self-energy can then show finite broadened levels shifted from stationary one-particle levels. An up to date linearly in time numerical method for solving closed equations of one-particle density matrices in the framework of NEGF in the time domain for coupled electron-boson dynamics is presented in KarlEtAl . It satisfies conservation laws despite being perturbative.
The route to kinetic equations we take with relaxator Liouville dynamics is close in spirit to the NEGF method, but the construction of the self-energy is formally exact and non-perturbative owing to a projector method. Our procedure has the advantage to immediately leading us to a closed equation for the one-particle density operator via a one-particle propagator. It makes the origin of irreversible behavior apparent by the relaxator appearing in this equation. The corresponding spectral shift and relaxator in our approach correspond to the self-energy in the NEGF approach.
@@ V Linear response for relaxator Liouville dynamics
Now, we expand this equation in powers of $\delta L{F}_{\tilde{\omega}}$ and estimate the regime of reliable linear response by requiring that the linear contribution to $\rho(z)$ is smaller than the contribution without the field. Since $(z-L(z))^{-1}$ scales with a characteristic system time, e.g. $t_{s}$ and the coupling Liouville scales with a characteristic rate as the inverse of a field induced transition time $t_{\rm F_{\omega}}$, the condition for the linear response theory to apply reads
$t_{s}\ll t_{\rm F_{\omega}}\,.$ (97)
@@ VI Weak coupling approximation
Here $N(\Omega)$ is the average equilibrium number of modes and $\nu(|\Omega|)$ the continuous density of frequency levels of bath modes. However, there is a problematic property of the linear coupling to independent bath modes: the correlation function $\gamma(\Omega)$ has no characteristic time scale involved besides the time scale inherent in $\rho_{\rm env}$ (e.g. the temperature time scale $T^{-1}$ for environmental equilibrium) and the inverse time scale inherent in the density of levels $\nu(|\Omega|)$ and the inverse time scale set by the coupling strength $\kappa(|\Omega|)$. The characteristic time scale of an environmental correlation time $\tau_{e}$ is missing as an ingredient in a bath of independent modes. Typically, for independent modes in the limit of large volume the density of levels becomes a power low $\sim\Omega^{p}$ with $p\geq 1$ such that $\gamma(0)\to 0$. However, for a realistic long time correlation function with decay on a scale $\tau_{e}$ one should find $\gamma(0)\propto\tau_{e}\not=0$. Therefore, one shouldn’t take the correlation function of (144) literally, but as a means of parametrization to be adjusted phenomenologically Weiss or to be improved by couplings which allow to introduce characteristic correlation time scales of the environment. Allowing for dependence of the coupling strength on the system’s energy levels already makes the linear coupling more reliable for certain aspects of the dynamics KarlEtAl . For a deeper discussion of limitations of the linear coupling to independent bath modes we refer to AlickiQDS and Horn2009 . In the following we will not rely on the linear coupling expression (144) but stick to the general properties of $\gamma(\Omega)$ for a bath in canonical equilibrium.
@@ VII Markov approximation: semi-group dynamics
as a semi-group-generator, because we want to be sure that its stationary limit is as close as possible with that of the original relaxator Liouville. The condition for the absence of memory (157) is quite restrictive and often not fulfilled in a realistic system when times between $\delta t_{s}$ and $\tau_{r}$ are addressed. One may ask why the semi-group approach that assumes the absence of memory (and absence of initial correlations) is still quite successful in describing several features of open systems. We think the answer can be given by looking at the solution (34) of the relaxator Liouville dynamics in its spectral representation. The solution (38) shows a superposition of damped oscillations which could as well be described as the solution of a frequency independent Liouville $L$ with its eigenfrequencies adopted to the effective eigenfrequencies $z_{k}$ (37) of the full $L(z)$. In particular, when fast oscillatory behavior is averaged out, the Markov approximation, still capturing relaxation, becomes better.
@@
VIII Summary and discussion
We developed the relaxator Liouville dynamics (Eqs. (4),(16) - (19)) for relevant degrees of freedom in a macroscopic system from a microscopic von Neumann equation for an isolated total system of all degrees of freedom. The irrelevant degrees of freedom act as an environment on the relevant degrees of freedom. The relaxator $\Gamma(\omega)$ indicates the onset of irreversibility. It can be expressed by a Hamiltonian modeling of system environment coupling (Eq. (70)). For a weak coupling the relaxator Liouville is determined by environmental correlation functions and hopping operators in the system (Eq. (122)). The relaxator Liouville dynamics has generically a unique stationary state (Eq. (48)) which is reached after a characteristic relaxation time that can be calculated from the relaxator Liouville’s effective eigenvalues (Eq. (39)). For equilibrium conditions the system reaches canonical thermal equilibrium with a temperature adopted to the environment.
@@ VIII Summary and discussion
For situations of current carrying stationary states in the presence of perturbing external fields we derived linear response expressions for dynamic susceptibilities (Eq. (102)). The derived susceptibilities coincide with Kubo’s susceptibilities in a regime where the energy eigenstates of the unperturbed system have undetectable large lifetimes and the probing frequency is much larger than the typical level spacing. Despite the practically infinite lifetimes of energy eigenstates, the response can show an collective irreversible character as the spectrum of the system becomes dense relative to the frequency. The corresponding time scale of irreversibility, a decay time of correlations within the system, can be extracted from the susceptibility (Eq. (117)). In the limit of separated system levels with finite life times the linear response susceptibility allows to extract lifetimes and level shifts (Eq. (106)).
@@
As the source of irreversibility we have identified time scale separations. For a system in contact with a surrounding bath the time to resolve the spectrum of the bath is much larger than the time to resolve the spectrum of the isolated system. For a macroscopic system in general the dynamics of the so-called relevant variables becomes irreversible when the time to resolve the spectrum of isolated relevant variables is much smaller than the time to resolve the spectrum of the total system. In transport situations with driving fields, irreversibility and dissipation of energy emerges when they are macroscopic or subject to macroscopic contacts such that the level spacing of the system cannot be resolved with the frequency of the applied field. In other words: the time to resolve the system’s spectrum is much larger as any time we can afford to measure the response of the system to applied fields.
Our general framework for modeling irreversible systems offers a generalization of the von Neumann dynamics for isolated relevant degrees of freedom to a relaxator Liouville dynamics of relevant degrees of freedom in open systems. It captures non-Markovian dynamics by the frequency dependence of $L(\omega)$. The total Hamiltonian is however assumed to be time independent (except for the investigation of linear response in Chap. V) and thus the total energy is conserved. Degrees of freedom are permanently treated as relevant or as irrelevant. Under these restrictions the full time dependence of environmental correlation functions can enter the construction of $L(\omega)$ and of $\rho_{0}(\omega)$ as given by Eqs. (16)-(19) and (69)-(74). The resolution of $L(\omega)$ is assumed to be reliable down to a frequency resolution of the typical level spacing of isolated relevant variables, $1/t_{s}$. Since this allows for non-Markovian dynamics including initial correlations with the environment a detailed time evolution of relevant degrees of freedom is possible. However this framework is aimed at and adopted to a situation where the number of irrelevant (environmental) degrees of freedom is huge compared to the number of relevant degrees of freedom. Due to the spectral condition ($t_{s}<<t_{e}$ (the level density of the total system cannot be resolved on the time scale $t_{s}$) the individual interaction processes between system and environment and within the environment are fast and the time evolution of such individual processes cannot be resolved in detail. If both $t_{s}$ and $t_{e}$ are of the same order the whole approach of a relaxator Liouville does no longer make sense. Both sets of variables (relevant and irrelevant) then have to be treated on the same level of resolution as can be done by the methods discussed in HoffEtAl . In HoffEtAl a cavity of two levels interacting with 400 photon modes is considered to describe precisely the emission, reflection, absorption and re-emission with interference of a one-photon wave packet. Such detailed interaction dynamics is beyond the scope of a Liouville relaxator modeling and other methods as investigated in HoffEtAl are surely superior. For $t_{e}\sim t_{s}$ is is hard to reach relaxation to steady state situations. The domain of applicability of a relaxator Liouville modeling is $t_{s}<<t_{e}$. It is adjusted for studying the non-Markovian approach to long time dynamics and steady state situations. With available spectral representation of $L(\omega)$ the full dynamics of relevant variables is given by (38). Whenever it comes to the irreversible time evolution of relevant degrees of freedom due to many individual interaction processes captured statistically by means of a relaxator $\Gamma(\omega)$ the relaxator Liouville modeling device is general and flexible. So far, the construction of $L(\omega)$ is straightforward in the weak coupling approximation where $\Gamma(\omega)$ separates into bath correlation functions and hopping operators in the system. To go beyond the weak coupling approximation by systematic methods that keep the total energy conserved is left for further investigations as well as applications to specific systems.
@@
Acknowledgements The authors acknowledge valuable comments during the process of writing from Victor Albert (2018), Heinz-Peter Breuer (2018) , John Chalker (2024), Rochus Klesse (2017), Dimitry Polyakov (2025), János Polonyi (2017, 2025) and Peter Wölfle (2025). The authors acknowledge support by the University of Cologne for providing open acces for publication in AdP.
Martin Janßen would like to express his deep gratitude for having shared many years with the late co-author János Hajdu. He was an instructive scholar and a wonderful person. As a scholar, he was distinguished by his comprehensive erudition and pedagogical skills, and as a person, by his great attentiveness and helpfulness.